OpenStudy (anonymous):

what is ^3 square root of 64

OpenStudy (anonymous):

you mean \[(\sqrt{64})^3\] or the cube root of 64?

OpenStudy (anonymous):

the cube root of 64 is 4

OpenStudy (anonymous):

cube root of 64

OpenStudy (anonymous):

2^2=sqrt(64) ===> it's 2

OpenStudy (anonymous):

o ok thanks!!

OpenStudy (anonymous):

\[\sqrt[3]{64}=4\]

OpenStudy (anonymous):

2^3=sqrt(64)

OpenStudy (anonymous):

your right kush

OpenStudy (amistre64):

3rt(64) is alot more readable

OpenStudy (anonymous):

remember 64 = 4*4*4 = 4^3 \[\sqrt[3]{4^3} =4\] since s=cube root itself means ^1/3 and 4^3^1/3 = 4^(3*1/3) = 4^1 = 4 since a^b^c=a^bc

OpenStudy (anonymous):

hope u must have understood these type of questions ,,,,,,, :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

its 4